# Hyper compressed fine-tuning of large foundation models with Quantum Inspired Adapters

### v1

@@ A.2.1 Compound Configurations
A.2.2 Orthogonality
To better understand the impact of keeping the adapter parameters orthogonal, we reran the experiments on STS-B but without cayley parameterization. The results are compared with their orthogonal counterpart in Figure  3.
@@ A.2.2 Orthogonality
A.2.3 Constructing adapters from alternate operations on minors
We also reran the experiments on STS-B with different operations on the minors as detailed in Section  4.3. The results are compiled in Figure 8.
@@ A.2.3 Constructing adapters from alternate operations on minors
A.2.4 Rank and Multi-adapter Analysis
@@
Table 6: Impact of Rank $r=\nicefrac{{d}}{{b}}$ and Number of Adapters on Accuracy and Parameter Count. The best performing configuration in absolute performance is in bold. The results are also visualised in Figure 9.
@@
Figure 9: Relationship Between Rank, $r:=\nicefrac{{d}}{{b}}$, Number of Adapters, and Accuracy Across Compound Patterns

### v2

@@ D.4 Ablation studies on STS-B dataset
To further understand the impact of different configuration setups, we run ablation studies on a dataset from the GLUE benchmark, specifically STS-B.
D.4.1 Compound Configurations
As illustrated in Figure 9, we explore how different configurations of QuIC Adapters perform on the STS-B dataset - an illustration of Table 4 in the main text. We note that the presence of $\mathcal{C}_{1}$ adapter with higher orders show the best performance while giving significant parameter reductions compared to only having higher order adapters ($\mathcal{C}_{2}$ or $\mathcal{C}_{3}$).
D.4.2 Orthogonality
Figure 10: Impact of enforcing orthogonality in adapters, for different compound configurations using STS-B.
To better understand the impact of keeping the adapter parameters orthogonal, we reran the experiments on STS-B but without cayley parameterization. The results are compared with their orthogonal counterpart in Figure  10.
@@ D.4.2 Orthogonality
D.4.3 Constructing adapters from alternate operations on minors
We also reran the experiments on STS-B with different operations on the minors as referred to in the main text. The results are compiled in Figure 11.
@@ D.4.3 Constructing adapters from alternate operations on minors
D.4.4 Rank and Multi-adapter Analysis
@@
Table 7: Impact of Rank $r=\nicefrac{{d}}{{b}}$ and number of adapters. The best performing configuration in absolute performance is in bold. The results are also visualized in Figure 12.
@@
Figure 12: Relationship between rank, $r:=\nicefrac{{d}}{{b}}$, number of adapters, and accuracy across compound configurations
@@
Appendix E Hyperparameters and Experimental Details
We report the key hyperparameters and experimental settings used across all benchmarks. All experiments were conducted on a single NVIDIA A80 GPU.
E.1 GLUE Benchmark
We evaluate on a subset of GLUE tasks: SST-2, CoLA, MRPC, and STS-B. Table 8 details the main hyperparameters for DeBERTaV3-base. All models are finetuned with AdamW optimizer and linear learning rate decay.
Table 8: Hyperparameters for GLUE (DeBERTaV3-base)
SST-2 CoLA MRPC STS-B
Batch Size 32 32 32 32
# Epochs 2 5 14 11
Learning Rate 2e-4 4e-4 9e-4 7e-4
Dropout 0.1 0.05 0.1 0.1
Max Sequence Length 128 64 320 128
E.2 VTAB-1K
We report results on five representative VTAB-1K tasks: CIFAR100, Pets, SVHN, Resisc45, and DMLab. All experiments use Adam optimizer and cosine learning rate schedule. The primary hyperparameter is the initial learning rate, set per task as in Table 9.
Table 9: Learning Rates for VTAB-1K Tasks
Dataset Learning Rate
CIFAR100 8e-4
Pets 3e-4
SVHN 3e-3
Resisc45 5e-4
DMLab 2e-3
E.3 MATH10K
For MATH10K experiments, we use a batch size of 4, AdamW optimizer, and a linear learning rate scheduler with an initial rate of 3e-4.
E.4 DROP
On DROP, we set batch size to 4, use AdamW optimizer, linear scheduler, and a learning rate of 1e-4.